Determine whether or not the following points are collinear. (3,1), (5,5), and (7,9)
Answer by stanbon(75887)

Determine whether or not the following points are collinear.
(3,1), (5,5), and (7,9)
---
Check the slope between each pair.
If the points are collinear the slopes should be equal.
---
Ex:
(5-1)/(5-3) = 4/2 = 2
(9-1)/(7-3) = 8/4 = 2
etc.
==============================
Cheers,
Stan H.